# From Restless to Contextual: A Thresholding Bandit Reformulation for Finite-horizon Improvement

### v4

@@ Abstract
This paper addresses the poor finite-horizon performance of existing online restless bandit (RB) algorithms, which stems from the prohibitive sample complexity of learning a full Markov decision process (MDP) for each agent. We argue that superior finite-horizon performance requires rapid convergence to a high-quality policy. Thus motivated, we introduce a reformulation of online RBs as a budgeted thresholding contextual bandit, which simplifies the learning problem by encoding long-term state transitions into a scalar reward. We prove the first non-asymptotic optimality of an oracle policy for a simplified finite-horizon setting. We propose a practical learning policy under a heterogeneous-agent, multi-state setting, and show that it achieves a sublinear regret, achieving faster convergence than existing methods. This directly translates to higher cumulative reward, as empirically validated by significant gains over state-of-the-art algorithms in large-scale heterogeneous environments. Our work provides a new pathway for achieving practical, sample-efficient learning in finite-horizon RBs.
@@ 1 Introduction
Restless bandits extend classic contextual bandits by modeling state transitions and incorporating per-step budget constraints. In a RB, a set of independent arms (agents) evolves over discrete time steps as MDP, with the goal of dynamically allocating a fixed number of active plays to maximize cumulative rewards. However, this modeling power comes with significant learning challenges. Even in the offline setting with known model parameters, RB is PSPACE-hard (Papadimitriou & Tsitsiklis, 1999). This computational intractability has directed much of the research toward policies that provide asymptotic guarantees, such as Whittle’s Index Policy (WIP) (Whittle, 1988; Weber & Weiss, 1990; Hsu, 2018; Kadota et al., 2016) and linear programming-based methods (Verloop, 2016; Zhang & Frazier, 2021; Brown & Zhang, 2023; Hong et al., 2024; Gast et al., 2024; Xiong et al., 2022a), which are proven to approach optimal as the number of agents approaches infinity.
@@
This complexity carries over into the online setting, where transition dynamics are unknown. Prior online RB algorithms largely seek to learn the optimal policy by first estimating the full underlying MDP for each agent. Consequently, online RB methods are largely inspired by, and mirror the standard frameworks of, reinforcement learning (RL). They predominantly fall into two canonical RL categories. Model-based approaches explicitly learn transition probabilities, typically within either episodic finite-horizon (Jung & Tewari, 2019; Wang et al., 2023b; Liang et al., 2024) or infinite-horizon (Ortner et al., 2012; Wang et al., 2020; Xiong et al., 2022b; Akbarzadeh & Mahajan, 2023; Xiong et al., 2022a) frameworks. Model-free methods focus on learning the Q-function specifically in the infinite-horizon setting (Killian et al., 2021; Xiong & Li, 2023; Xiong et al., 2024; Kakarapalli et al., 2024). A common strategy in both categories is to then integrate these learned models with WIP for decision-making. [^1] We note that while LP-based methods are powerful in the offline setting, they are generally more difficult to adapt for online learning. These online methods are designed to converge to their asymptotically optimal offline counterparts, where the convergence rate is measured by the notation of regret (formalized in Sec. 2.2).
@@
This reliance on standard RL frameworks means that the performance guarantees for these online algorithms are inherently asymptotic in the number of episodes or time steps. However, this focus constitutes a fundamental limitation in practical applications, [^2] Table A.1 provides key application areas for restless bandits. where horizons are finite due to budget cycles, program durations, or policy deadlines (del Río et al., 2024a, b). In domains like targeted patient interventions (Villar et al., 2015; Biswas et al., 2021; Mate et al., 2021) or resource allocation (Borkar et al., 2017; Brown & Smith, 2020), systems evolve continuously and cannot be reset, making the standard episodic RL framework impractical. Consequently, an algorithm’s performance is judged by its cumulative reward within a single, limited-time episode. An asymptotically optimal policy that requires a long time to learn offers no practical value if it fails to achieve a high reward within the operational horizon. This critical limitation of prior work motivates a fundamental shift away from the objective of learning complex MDPs, and towards a reformulation designed for sample-efficient learning in the non-episodic, finite-horizon setting. Our contributions are fourfold:
@@
  • A Novel Reformulation: we interpret regret as convergence rate (Sec. 2.2) and show that online algorithms may achieve reward maximization more effectively by targeting alternative objectives, which leads to accelerated convergence (Sec. 2.3). This motivates us to introduce a budgeted thresholding contextual multi-armed bandit (BT-CMAB), which simplifies the learning objective of RBs to a contextual bandit problem through carefully designed rewards (Sec. 2.4), enabling sample-efficient learning in finite horizons.
@@
  • First Non-asymptotic Finite-Horizon Optimality Guarantee: we establish the first finite-horizon non-asymptotic (valid for arbitrary number of agents) optimality guarantee for a tractable RB setting, proving our oracle greedy policy is optimal for a homogeneous 2-state model (Theorem 3.1). This provides a theoretical foundation absent in prior work.
@@ 1.1 Additional Related Work
Thresholding Bandit/MDP   Our paper relates to regret-minimizing thresholding bandit problems, where regret is defined by the gap between the arm’s mean and a predefined threshold when the mean falls below the threshold. Under this setting, minimax optimal (max is taken over all possible bandit instances and min taken over all possible learning algorithms) constant regret is achievable (Tamatsukuri & Takahashi, 2019; Michel et al., 2022; Feng et al., 2024). Our problem setup differs from the traditional thresholding bandit framework by incorporating context, where the arrival of contexts follows an MDP. In contrast, Hajiabolhassan & Ortner (2023) study threshold MDPs under an infinite horizon setting, seeking policies with average rewards exceeding a predefined threshold. We, however, focus on the finite-horizon setting.
@@
MAB Experimental Deisgn   Our incremental reward explicitly incorporates the trade-off between reward maximization and treatment effect estimation, previously studied in MAB with Pareto-optimal algorithms provided (Simchi-Levi & Wang, 2025; Zhong et al., 2021; Zhang & Wang, 2024). Unlike prior work that treats these aspects separately, ours inherently combines both aspects through reward that depends on intervention and non-intervention outcomes.
@@ 2 Finite-horizon Restless Bandits and Our Reformulation
We introduce restless bandits in Sec. 2.1 and reinterpret regret as convergence in Sec. 2.2. Sec. 2.3 argues that better finite-sample performance requires new design objectives, while maintaining reward-based evaluation. We then present our reformulation in Sec. 2.4.
@@ 2.1 Restless Bandits
In a finite-horizon RB problem, the learner is provided with $M$ independent arms, each representing one independent agent. Each agent $m$ is modeled as an MDP represented by the tuple $\left(\mathcal{S},\mathcal{A},P_{a}^{m},R^{m}\right)$, where $P_{a}^{m}(s,s^{\prime}):\mathcal{S}\times\mathcal{S}\to[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ under action $a$ for agent $m$, and $R^{m}(s):\mathcal{S}\to\mathbb{R}$ represents the expected reward that agent $m$ receives at state $s$. We focus on binary action, i.e., $\mathcal{A}=\{0,1\}$. We highlight that in RB, the expected reward, $R^{m}(s)$, is independent of the action, and the learner observes the reward immediately after arriving at the state.
@@
We permit non-stationarity in the rewards, provided that Assump. 2.1 is satisfied, offering a relaxation of the standard stationarity assumption typically imposed in literature.
The goal of the learner is to find the policy $\bm{\pi}:=\{\pi_{t}\}_{t\in[T]}$ that maximizes the expected cumulative reward, where the expectation is further taken over the randomized policy ($\bm{a}_{t}\sim\pi_{t}$), in addition to the states ($s_{t+1}^{m}\sim P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})$), subject to the full budget utilization constraint: [^3] Our model incorporates the standard full budget utilization constraint, consistent with the existing literature (Whittle, 1988; Zhang & Frazier, 2021; Hong et al., 2024).
@@
We will use $\bm{\pi}^{*}$ to denote the optimal policy corresponding to Problem (1). Since $\bm{\pi}^{*}$ is intractable and unknown (Papadimitriou & Tsitsiklis, 1999), WIP (Whittle, 1988) and LP-based policies (Hong et al., 2024; Xiong et al., 2022a) are proposed as approximations. These policies are proven to be asymptotically optimal as $M\to\infty$ for infinite-horizon rewards with homogeneous agents, a restrictive setting that often does not occur in practice.
@@ 2.2 Regret As Convergence Rate
$-\mathbb{E}_{\bm{\omega}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}l^{m}(s_{t}^{m},a_{t+1}^{m})\mid\bm{s}_{0}\right].$
Here, $l^{m}(s_{t}^{m},a_{t+1}^{m})$ serves as an instantaneous performance metric. Def. 2.2 uses the cumulative difference in this loss to measure the convergence rate of the online algorithm $\bm{\omega}$ to the reference policy $\bm{\pi}^{e}$. This intrinsically links the choice of the loss function $l^{m}$ to the objective of $\bm{\pi}^{e}$. For instance, in MAB, if $\bm{\pi}^{e}$ targets reward maximization, $l$ is immediate reward (Auer et al., 2002); if $\bm{\pi}^{e}$ targets thresholding, $l$ penalizes action on arms below threshold (Feng et al., 2024). Thus, $l$ always measures how well online algorithms are achieving the corresponding objectives of $\bm{\pi}^{e}$.
@@ 2.3 The Need for Indirect Reward Maximization in Online RB
Building on our reinterpretation of regret as convergence rate, we reconsider Problem (1)’s reward maximization objective. In finite-horizon settings, algorithms targeting alternative objectives with faster convergence to high-quality reference policies may achieve superior empirical performance. This approach addresses the fundamental limitation of direct methods for online RB—including those learning $\bm{\pi}^{*}$, WIP (Jung & Tewari, 2019; Wang et al., 2023b), or LP-based policies (Xiong et al., 2022a)—which typically suffer from slow convergence in practice. [^5] Although Xiong et al. (2022a) claim sublinear convergence for finite-horizon RB, their analysis relies on a generative model that permits restarting the MDP from arbitrary states—an assumption rarely satisfied in practice. Without this, their algorithm would incur linear convergence.
@@
Given these fundamental constraints on estimating transition probabilities, we pursue an alternative approach: selecting a reference policy that enables faster convergence by circumventing the need for direct transition probability estimation. We develop this in Sec. 2.4, focusing on reference policies that avoid this bottleneck while maintaining strong empirical performance.
@@ 2.4 BT-CMAB Reformulation
    Contextual bandit framework: Transition probabilities need not be explicitly estimated. It embeds nonstationarity into rewards, yielding policies that depend only on the current context. This reduces sample complexity when compared with MDPs (Simchi-Levi & Xu, 2022).
@@
$\phi^{m}(s_{t}^{m},a_{t+1}^{m}):$ $=\sum_{s_{t+1}^{m}\in\mathcal{S}}P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})R^{m}\left(s_{t+1}^{m}\right)$
@@
As in online RBs, at each time $t$, the learner observes $\bm{s}_{t}$ and the corresponding reward $\bm{r}_{t}$. The learner then updates the estimated uncertainty associated with the estimated $\phi^{m}(s_{t-1}^{m},a_{t}^{m})\;\forall m\in[M]$ according to Eq. (2), and uses it to guide the action $\bm{a}_{t+1}$, with $\sum_{m\in[M]}a_{t+1}^{m}=B$. The goal of the learner is to learn a policy that identifies and selects $B$ agents whose expected gain from taking the action (i.e., $a_{t+1}^{m}=1$) exceeds a predefined threshold, $\gamma$, in their current state $s_{t}^{m}$. To quantify this expected gain in state $s\in\mathcal{S}$, we define $I^{m}(s):=\phi^{m}(s,1)-\phi^{m}(s,0)$, representing the incremental reward an agent receives by taking the action versus not at state $s$. At time $t$, we say that an agent $m$ is in a good state if $I^{m}(s_{t}^{m})\geq\gamma$.
At each time $t$, let $G_{t}:=\{m\in[M]:I^{m}(s_{t}^{m})\geq\gamma\}$ denote the set of agents currently in a good state. In this budgeted thresholding contextual bandit formulation, the optimal policy, $\bm{\pi}^{b}$, involves randomly selecting $B$ agents from $G_{t}$ for intervention when $|G_{t}|\geq B$ and selecting $B$ agents according to the ordering of incremental reward otherwise.
@@
$c_{t}^{m}\left(a_{t+1}^{m}\right):=\left(\gamma-I^{m}(s_{t}^{m})\right)\cdot\mathbbm{1}\left\{I^{m}(s_{t}^{m})<\gamma\right\}\cdot\mathbbm{1}\left\{a_{t+1}^{m}=1\right\}.$ (3)

### v5

@@ Abstract
This paper addresses the poor finite-horizon performance of existing online restless bandit (RB) algorithms, which stems from the prohibitive sample complexity of learning a full Markov decision process (MDP) for each agent. We argue that superior finite-horizon performance requires rapid convergence to a high-quality policy. Thus motivated, we introduce a reformulation of online RBs as a budgeted thresholding contextual bandit, which simplifies the learning problem by encoding long-term state transitions into a scalar reward. We prove the first non-asymptotic optimality of an oracle policy for a simplified finite-horizon setting. We propose a practical learning policy under a heterogeneous-agent, multi-state setting, and show that it achieves a sublinear regret, achieving faster convergence than existing methods. This directly translates to higher cumulative reward, as empirically validated by significant gains over state-of-the-art algorithms in large-scale heterogeneous environments. The code is provided in github. Our work provides a new pathway for achieving practical, sample-efficient learning in finite-horizon RBs.
@@ 1 Introduction
Restless bandits extend classic contextual bandits by modeling state transitions and incorporating per-step budget constraints. In a RB, a set of independent arms (agents) evolves over discrete time steps as MDP, with the goal of dynamically allocating a fixed number of active plays to maximize cumulative rewards. However, this modeling power comes with significant learning challenges. Even in the offline setting with known model parameters, RB is PSPACE-hard (Papadimitriou and Tsitsiklis, 1999). This computational intractability has directed much of the research toward policies that provide asymptotic guarantees, such as Whittle’s Index Policy (WIP) (Whittle, 1988; Weber and Weiss, 1990; Hsu, 2018; Kadota et al., 2016) and linear programming-based methods (Verloop, 2016; Zhang and Frazier, 2021; Brown and Zhang, 2023; Hong et al., 2024; Gast et al., 2024; Xiong et al., 2022a), which are proven to approach optimal as the number of agents approaches infinity.
This complexity carries over into the online setting, where transition dynamics are unknown. Prior online RB algorithms largely seek to learn the optimal policy by first estimating the full underlying MDP for each agent. Consequently, online RB methods are largely inspired by, and mirror the standard frameworks of, reinforcement learning (RL).
@@
They predominantly fall into two canonical RL categories. Model-based approaches explicitly learn transition probabilities, typically within either episodic finite-horizon (Jung and Tewari, 2019; Wang et al., 2023b; Liang et al., 2024) or infinite-horizon (Ortner et al., 2012; Wang et al., 2020; Xiong et al., 2022b; Akbarzadeh and Mahajan, 2023; Xiong et al., 2022a) frameworks. Model-free methods focus on learning the Q-function specifically in the infinite-horizon setting (Killian et al., 2021; Xiong and Li, 2023; Xiong et al., 2024; Kakarapalli et al., 2024). A common strategy in both categories is to then integrate these learned models with WIP for decision-making. [^1] We note that while LP-based methods are powerful in the offline setting, they are generally more difficult to adapt for online learning. These online methods are designed to converge to their asymptotically optimal offline counterparts when agents are homogeneous, where the convergence rate is measured by the notion of regret (formalized in Sec. 2.2).
@@
This reliance on standard RL frameworks means that the performance guarantees for these online algorithms are inherently asymptotic in the number of episodes or time steps. However, this focus constitutes a fundamental limitation in practical applications, [^2] Table A.1 provides key application areas for RBs. where horizons are finite due to budget cycles, program durations, or policy deadlines (delrío2024; delrío2024Pharmacy). In domains like targeted patient interventions (Villar et al., 2015; Biswas et al., 2021; Mate et al., 2021) or resource allocation (Borkar et al., 2017; Brown and Smith, 2020), systems evolve continuously and cannot be reset, making the standard episodic RL framework impractical.
Consequently, an algorithm’s performance is judged by its cumulative reward within a single, limited-time episode. An asymptotically optimal policy that requires a long time to learn offers no practical value if it fails to achieve a high reward within the operational horizon. This critical limitation of prior work motivates a fundamental shift away from the objective of learning complex MDPs, and towards a reformulation designed for sample-efficient learning in the non-episodic, finite-horizon setting. Our contributions are fourfold:
@@
  • A Novel Reformulation: we interpret regret as convergence rate (Sec. 2.2) and show that online algorithms may achieve higher cumulative reward in a finite horizon by targeting alternative design objectives, which leads to accelerated convergence (Sec. 2.3). This motivates us to introduce a budgeted thresholding contextual multi-armed bandit (BT-CMAB), which simplifies the learning objective of RBs to a contextual bandit problem through carefully designed rewards (Sec. 2.4), enabling sample-efficient learning in finite horizons.
@@
  • First Non-asymptotic Finite-Horizon Optimality Guarantee: we establish the first finite-horizon non-asymptotic (valid for an arbitrary number of agents) optimality guarantee for a tractable RB setting, proving our oracle greedy policy is optimal for a homogeneous 2-state model (Theorem 3.1). This provides a theoretical foundation absent in prior work.
@@ 1.1 Additional Related Work
Thresholding Bandit/MDP   Our paper relates to regret-minimizing thresholding bandit problems, where regret is defined by the gap between the arm’s mean and a predefined threshold when the mean falls below the threshold. Under this setting, minimax optimal (max is taken over all possible bandit instances and min taken over all possible learning algorithms) constant regret is achievable (Tamatsukuri and Takahashi, 2019; Michel et al., 2022; Feng et al., 2024). Our problem setup differs from the traditional thresholding bandit framework by incorporating context, where the arrival of contexts follows an MDP. In contrast, Hajiabolhassan and Ortner (2023) study threshold MDPs under an infinite horizon setting, seeking policies with average rewards exceeding a predefined threshold. We, however, focus on the finite-horizon setting.
@@
MAB Experimental Design   Our incremental reward explicitly incorporates the trade-off between reward maximization and treatment effect estimation (i.e., estimating the difference in rewards under $A=0$ and $A=1$), previously studied in MAB with Pareto-optimal algorithms provided (Simchi-Levi and Wang, 2025; Zhong et al., 2021; Zhang and Wang, 2024). Unlike prior work that treats these aspects separately, ours inherently combines both aspects through reward that depends on intervention (i.e., $A=1$) and non-intervention (i.e., $A=0$) outcomes.
@@ 2 Finite-horizon Restless Bandits and Our Reformulation
We introduce restless bandits in Sec. 2.1 and reinterpret regret as convergence in Sec. 2.2. Sec. 2.3 argues that better finite-sample performance requires new design objectives, while maintaining reward-based evaluation. We present our reformulation in Sec. 2.4.
@@ 2.1 Restless Bandits
In a finite-horizon RB problem, the learner is provided with $M$ independent arms, each representing one independent agent. Each agent $m$ is modeled as an MDP represented by the tuple $\left(\mathcal{S},\mathcal{A},P_{a}^{m},R^{m}\right)$, where $P_{a}^{m}(s,s^{\prime}):\mathcal{S}\times\mathcal{S}\to[0,1]$ represents the transition probability from state $s$ to $s^{\prime}$ under action $a$ for agent $m$, and $R^{m}(s):\mathcal{S}\to\mathbb{R}$ represents the expected reward that agent $m$ receives at state $s$. We focus on binary action, i.e., $\mathcal{A}=\{0,1\}$. We highlight that in classical RBs, the expected reward, $R^{m}(s)$, is independent of the action, and the learner observes the reward immediately after arriving at the state.
@@
We permit non-stationarity in the rewards, provided that Assumption 2.1 is satisfied, offering a relaxation of the standard stationarity assumption typically imposed in literature.
The goal of the learner is to find the policy $\bm{\pi}:=\{\pi_{t}\}_{t\in[T]}$ that maximizes the expected cumulative reward, where the expectation is further taken over the randomized policy ($\bm{a}_{t}\sim\pi_{t}$), in addition to the states ($s_{t+1}^{m}\sim P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})$), subject to the full budget utilization constraint: [^3] Our model incorporates the standard full budget utilization constraint, consistent with the existing literature (Whittle, 1988; Zhang and Frazier, 2021; Hong et al., 2024).
@@
We will use $\bm{\pi}^{*}$ to denote the optimal policy corresponding to Problem (1). Since $\bm{\pi}^{*}$ is intractable and unknown (Papadimitriou and Tsitsiklis, 1999), WIP (Whittle, 1988) and LP-based policies (Hong et al., 2024; Xiong et al., 2022a) are proposed as approximations. These policies are proven to be asymptotically optimal as $M\to\infty$ for infinite-horizon rewards with homogeneous agents, a restrictive setting that often does not occur in practice.
@@ 2.2 Regret As Convergence Rate
$-\mathbb{E}_{\bm{\omega}}\left[\sum_{m=1}^{M}\sum_{t=0}^{T}l^{m}(s_{t}^{m},a_{t+1}^{m})\mid\bm{s}_{0}\right],$
where $l^{m}(s_{t}^{m},a_{t+1}^{m})$ serves as an instantaneous performance metric where the choice of $l^{m}$ depends on the objective of $\bm{\pi}^{e}$, which we elaborate below. For heterogeneous agents with single-episode finite horizons, existing RB oracle policies are not optimal and hence not comparable. Thus, the regret $\mathcal{R}_{l}^{\bm{\pi}^{e}}(\bm{\omega})$ measures the convergence rate of the learning algorithm $\omega$ to its corresponding oracle (or reference) policy $\bm{\pi}^{e}$, where sublinear regret implies that $\omega$ converges to $\bm{\pi}^{e}$ within sublinear time.
The choice of the loss function $l^{m}$ is directly linked to the objective of $\bm{\pi}^{e}$. For instance, in MAB, if $\bm{\pi}^{e}$ targets reward maximization, $l$ is immediate reward (Auer et al., 2002); if $\bm{\pi}^{e}$ targets thresholding, $l$ penalizes action on arms below threshold (Feng et al., 2024). Thus, $l$ always measures how well online algorithms are achieving the corresponding objectives of $\bm{\pi}^{e}$.
@@ 2.3 The Need for Indirect Reward Maximization in Online RB
Building on our reinterpretation of regret as convergence rate, we reconsider Problem (1)’s reward maximization objective. In finite-horizon settings, algorithms targeting alternative objectives with faster convergence to high-quality reference policies may achieve superior empirical performance. This approach addresses the fundamental limitation of direct methods for online RB—including those learning $\bm{\pi}^{*}$, WIP (Jung and Tewari, 2019; Wang et al., 2023b), or LP-based policies (Xiong et al., 2022a)—which typically suffer from slow convergence in practice. [^5] Although Xiong et al. (2022a) claim sublinear convergence for finite-horizon RB, their analysis relies on a generative model that permits restarting the MDP from arbitrary states—an assumption rarely satisfied in practice. Without this, their algorithm would incur linear convergence.
@@
Given these fundamental constraints on estimating transition probabilities, we pursue an alternative approach: proposing an oracle (or reference) policy that enables faster convergence by circumventing the need for direct transition probability estimation. We develop this in Sec. 2.4, focusing on reference policies that avoid this bottleneck while maintaining strong empirical performance.
@@ 2.4 BT-CMAB Reformulation
    Contextual bandit framework: Transition probabilities need not be explicitly estimated. It embeds nonstationarity into rewards, yielding policies that depend only on the current context. This reduces sample complexity when compared with MDPs (Simchi-Levi and Xu, 2022).
@@
$\phi^{m}(s_{t}^{m},a_{t+1}^{m})$ $\coloneq\sum_{s_{t+1}^{m}\in\mathcal{S}}P^{m}_{a_{t+1}^{m}}(s_{t}^{m},s_{t+1}^{m})R^{m}\left(s_{t+1}^{m}\right)$
@@
As in online RBs, at each time $t$, the learner observes $\bm{s}_{t}$ and the corresponding reward $\bm{r}_{t}$. The learner then updates the estimated uncertainty associated with the estimated $\phi^{m}(s_{t-1}^{m},a_{t}^{m})\;\forall m\in[M]$ according to Eq. (2), and uses it to guide the action $\bm{a}_{t+1}$, with $\sum_{m\in[M]}a_{t+1}^{m}=B$. The goal of the learner is to learn a policy that identifies and selects $B$ agents whose expected gain from taking the action (i.e., $a_{t+1}^{m}=1$) exceeds a predefined threshold, $\gamma$, in their current state $s_{t}^{m}$. To quantify this expected gain in state $s\in\mathcal{S}$, we define
$I^{m}(s):=\phi^{m}(s,1)-\phi^{m}(s,0),$
representing the incremental reward an agent receives by taking the action versus not at state $s$. At time $t$, we say that an agent $m$ is in a good state if $I^{m}(s_{t}^{m})\geq\gamma$.
At each time $t$, let $G_{t}:=\{m\in[M]:I^{m}(s_{t}^{m})\geq\gamma\}$ denote the set of agents currently in a good state. In this BT-CMAB formulation, the optimal policy, $\bm{\pi}^{b}$, involves randomly selecting $B$ agents from $G_{t}$ for intervention when $|G_{t}|\geq B$ and selecting $B$ agents according to the ordering of incremental reward otherwise.
@@
$c_{t}^{m}\left(a_{t+1}^{m}\right):=\left(\gamma-I^{m}(s_{t}^{m})\right)\mathbb{I}\left\{I^{m}(s_{t}^{m})<\gamma,a_{t+1}^{m}=1\right\}.$ (3)